OpenStudy (anonymous):

simplify the rational expression 4x-32 ------ 4x+28

OpenStudy (solomonzelman):

H I N T : \[4x-32=4(x-8)\]\[4x+28=4(x+7)\]

OpenStudy (anonymous):

i dont know

OpenStudy (solomonzelman):

\[\frac{4(x-8)}{4(x+7)}\] \[\frac{\cancel{4} (x-8)}{\cancel{ 4} (x+7)}\]do you know now ?

OpenStudy (anonymous):

yes thank you